# Adaptive Rank Allocation for Federated Parameter- Efficient Fine-Tuning of Language Models

### v3

@@ Abstract
Pre-trained Language Models (PLMs) have demonstrated their superiority and versatility in modern Natural Language Processing (NLP), effectively adapting to various downstream tasks through further fine-tuning. Federated Parameter-Efficient Fine-Tuning (FedPEFT) has emerged as a promising solution to address privacy and efficiency challenges in distributed training for PLMs on resource-constrained local devices. However, our measurements reveal two key limitations of FedPEFT: heterogeneous data across devices exacerbates performance degradation of low-rank adaptation, and a fixed parameter configuration results in communication inefficiency. To overcome these limitations, we propose FedARA, a novel Adaptive Rank Allocation framework for federated parameter-efficient fine-tuning of language models. Specifically, FedARA employs truncated Singular Value Decomposition (SVD) adaptation to enhance similar feature representation across clients, significantly mitigating the adverse effects of data heterogeneity. Subsequently, it utilizes dynamic rank allocation to progressively identify critical ranks, effectively improving communication efficiency. Lastly, it leverages rank-based module pruning to automatically remove inactive modules, steadily reducing local computational cost and memory usage in each federated learning round. Extensive experiments show that FedARA consistently outperforms baselines by an average of 6.95% to 8.49% across various datasets and models under heterogeneous data while significantly improving communication efficiency by 2.40$\times$. Moreover, experiments on various edge devices demonstrate substantial decreases in total training time and energy consumption by up to 48.90% and 46.95%, respectively.
Index Terms:
Federated Learning, Communication Efficiency, Data Heterogeneity, Parameter-Efficient Fine-Tuning.
@@ I Introduction
Although PLMs have demonstrated superior performance on general-purpose tasks [5, 6], they still require fine-tuning with local data to achieve optimal results in specific domains. However, since this data is typically generated by users, directly collecting it poses a significant risk of privacy leakage. To address this issue, federated learning (FL) is considered an effective approach [7], as it protects data privacy by keeping raw data on local devices, ensuring compliance with regulations such as GDPR [8] and the Data Protection Act [9]. The FL process involves a central server orchestrating the aggregation of local parameter updates from individual clients, with this cycle repeating until training converges.
@@
Figure 1: The complete training process of Federated Parameter-Efficient Fine-Tuning (FedPEFT), where local updates utilize Low-Rank Adaptation (LoRA) while keeping the pre-trained model frozen.
@@
However, the massive scale of PLMs, often with hundreds of millions to billions of parameters [10], incurs substantial computation and communication costs in FL, rendering deployment on resource-constrained devices nearly impossible. To address this issue, a promising solution is to adopt Parameter-Efficient Fine-Tuning (PEFT) methods, which train only a minimal set of parameters using adaptable modules while keeping the pre-trained model frozen [11, 12, 13, 14, 15]. Among these PEFT methods, Low-Rank Adaptation (LoRA) [15] has gained widespread adoption for its effectiveness and efficiency, achieved by introducing two trainable low-rank matrices alongside the pre-trained model for diverse task adaptation. By combining PEFT with FL, distributed data resources on local devices can be fully leveraged for collaborative fine-tuning in a privacy-preserving manner. Consequently, Federated Parameter-Efficient Fine-Tuning (FedPEFT) shows significant advancements and finds wide-ranging applications across various domains [16, 17, 18, 19, 20, 21, 22]. Building on this progress, FedPEFT has increasingly gained recognition as a key method for distributed fine-tuning on local devices [23, 24]. Figure 1 shows the complete FedPEFT training process with LoRA in local clients.
@@
Nevertheless, FedPEFT still faces two major challenges when applied to local devices. 1) FedPEFT exacerbates the impact of non-independent and identically distributed (non-IID) data on model performance. Although performance degradation caused by non-IID data is a well-known challenge in FL, our preliminary measurements show that FedPEFT exacerbates this issue significantly. Specifically, on the 20News [25] dataset using the DistilBERT [2] model, FedPEFT methods exhibit an average accuracy drop of 14.62% under non-IID conditions compared to IID settings—more than double the 6% drop observed with federated full fine-tuning [19]. 2) Fixed rank configurations in FedPEFT incur dual losses in performance and communication cost. Low-rank parameters can be inserted into various positions within the model, such as layers or components like the query, key, and value matrices. However, static allocations often underfit high-sensitivity positions, leading to suboptimal performance—our experiments reveal accuracy gaps of up to 9.76% across different configurations. Meanwhile, redundant parameters assigned to less critical components are repeatedly transmitted during FL rounds without corresponding performance gains, resulting in communication inefficiency.
@@
These dual losses become more severe under non-IID settings, where data heterogeneity intensifies the subspace divergence of federated low-rank adaptation among client-specific models. This divergence leads to greater variability in rank requirements, making fixed rank configurations increasingly misaligned with actual capacity needs. Consequently, both performance degradation and communication overhead are further amplified. Therefore, these two challenges—data heterogeneity and fixed rank allocation—must be jointly addressed to fully unlock the potential of FedPEFT in practical federated learning scenarios.
@@
Results. We implemented FedARA and evaluated it on three representative edge devices, NVIDIA Jetson AGX Orin, Orin Nano, and Raspberry Pi 5, whose hardware profiles reflect typical resource-constrained settings. We utilize three pre-trained models across five diverse NLP datasets. FedARA outperforms baselines [22, 26, 19, 20] on classification tasks by an average of 6.95% to 8.49% under non-IID data while significantly improving communication efficiency by an average of 2.40$\times$. Remarkably, the total training time on the above devices also decrease by 27.65%, 46.57%, and 48.90%, respectively, and the energy consumption drops by 46.95% on Orin Nano. Ablation experiments reveal that truncated SVD adaptation improves the average accuracy by 7.71%, dynamic rank allocation reduces communication overhead by 70.82% per round while maintaining on-par performance, and rank-based module pruning lowers the average local computation and peak GPU memory usage per round by 10.81% and 31.67%, respectively.
@@
  • We validate the effectiveness of FedARA through extensive experiments across various datasets and models on advanced edge devices, demonstrating that FedARA consistently improves performance under non-IID data and significantly boosts systems efficiency in both communication and computation.
@@ II-A Federated Parameter-Efficient Fine-Tuning (FedPEFT)
FedPEFT orchestrates PEFT strategies within the FL framework to efficiently update local models, drastically reducing communication overhead by only transmitting a small number of PEFT modules. These PEFT modules can be replaced by Adapter [11, 12], Prompt[13, 14], or LoRA [15]. Among them, LoRA-based methods are especially popular because they introduce no additional latency during inference. FedPEFT not only preserves privacy but also effectively aggregates diverse knowledge across clients, facilitating the collaborative development of a high-quality global model. By harnessing the advantages of both PEFT and FL, FedPEFT has become an essential technique for deploying fine-tuning on resource-constrained devices [23, 24, 20, 18, 17, 21].
As illustrated in Figure 1, FedPEFT training process unfolds as follows: Step 1: A cloud server selects a subset of clients to broadcast the initial parameters. Step 2: On these clients, the weights of the pre-trained model are frozen, and trainable PEFT modules are inserted for local training. Step 3: Only these PEFT modules are updated to the cloud server from each client. Step 4: The cloud server aggregates these updates from different clients. This cycle (Steps 1-4) is repeated for multiple rounds in FL until the model converges.
@@ II-B Preliminary Measurements
Observation-1: FedPEFT is more effective than FedFFT on advanced edge devices. FedPEFT comprises two main methods: adapter-based [23] and LoRA-based methods [16, 17, 18, 19, 21]. LoRA-based methods have gradually replaced adapter-based methods to avoid additional inference latency. According to established work [23], FedPEFT achieves comparable accuracy to federated full fine-tuning (FedFFT) while significantly reducing local training time by 1.6$\times$, communication cost by 200$\times$, and memory usage by 1.5$\times$ on previous-generation edge devices like NVIDIA Jetson TX2, Nano, and Raspberry Pi 4B. However, its impact on overall training time on advanced devices remains underexplored. To fill this gap, we evaluate FedLoRA on the 20News [25] dataset with the BERT [3] model using latest edge devices. Following FedPEFT’s typical settings [23], we set the communication bandwidth to 1MB/s. As depicted in Figure 2a, the results indicate that FedFFT’s overall training time is 42$\times$, 40$\times$, and 18$\times$ longer than FedLoRA on AGX Orin, Orin Nano, and RPi 5, respectively. FedPEFT demonstrates its effectiveness compared to FedFFT for cutting-edge embedded devices.
@@
Observation-2: The performance of FedPEFT significantly decreases under non-IID data. Figure 2b illustrates the accuracy of different FedPEFT methods using the DistilBERT [2] model for both IID (white bars) and non-IID data (black bars). The non-IID division follows the FedAvg setting [7], with a severe class imbalance in data distribution. The parameter sizes for Adapter-h (A-h) [11] and Adapter-p (A-p) [12] are set to match those of the LoRA method. Our measurements indicate a significant performance drop in all FedPEFT methods under non-IID conditions compared to IID conditions, with average accuracy reductions of 14.62% and 12.81% on the 20News [25] and News Category [27] datasets, respectively. Existing research lacks effective strategies to mitigate the adverse impact of non-IID data on FedPEFT performance.
@@
Observation-3: The contribution of PEFT modules in FedPEFT to performance varies their inserted positions, leading to suboptimal performance and inefficient communication in fixed parameter configurations. In FedPEFT, low-rank modules can be placed at various layers and components, such as Query (Q), Key (K), Value (V), Out (O), F1, and F2. Here, O denotes the linear layer in the multi-head attention block, while F1 and F2 represent the two linear layers of the feed-forward network. As shown in Figure 2c, we allocated a fixed rank budget to specific components or layers to evaluate their impact on accuracy. The left sub-figure shows the allocation of the entire rank budget to a single layer, while the right sub-figure illustrates its distribution across a single component in all layers. The findings reveal that deeper layers near the output often have a more significant effect than shallower ones, with components F1 and F2 typically exerting a greater influence than component O. Moreover, the impact of these inserted positions varies across datasets, suggesting that there is no one-size-fits-all configuration. Fixed rank configurations can result in inefficient communication, as they repeatedly transmit less impactful parameters over hundreds of rounds, yielding minimal performance gains under a limited parameter budget.
@@
Observation-4: Bottlenecks differ across different types of advanced edge devices. In experiments with LoRA-based methods [16, 17] on the News Category [27] dataset, given a communication bandwidth of 1MB/s, we found that the bottlenecks varied depending on the device type, as shown in Figure 2d. On high-end devices, communication is the predominant bottleneck, with communication-to-computation ratios of 4.44 and 5.35 for different models. In contrast, on Raspberry Pi devices, computation becomes the primary bottleneck, with communication-to-computation ratios of 0.67 and 0.72. Importantly, the bottleneck shifts towards communication as model parameters increase, consistent with recent studies [24].
@@
Insights: These observations underscore key directions for advancing our research. First, given its established advantages, such as improved training efficiency, future research could benefit from moving beyond comparisons with FedFFT and instead focusing on addressing the unique challenges it presents in FL. Second, the performance degradation of different PEFT methods under non-IID remains a critical barrier, necessitating further investigation into effective approaches. Third, fixed parameter configurations lead to communication inefficiency by repeatedly transmitting less impactful parameters over multiple FL rounds, particularly in resource-constrained environments. Finally, optimizing FedPEFT for edge devices requires further refinement, as different hardware presents distinct bottlenecks. To bridge these gaps, enhancing fine-tuning structures and dynamically adjusting parameters could be essential while simultaneously minimizing both communication and computational overhead.
@@ III Overview of FedARA
Algorithm 1 outlines the process of federated adaptive rank allocation. It starts with parameter initialization and client selection (line 5-7). The server prunes parameters using global rank masks (lines 8-9) and broadcasts them to the clients (line 10). Clients perform parallel local training (lines 11-12), updating parameters based on local datasets (lines 24). Local rank masks are generated from updated parameters (lines 25-26), determining the number of trainable parameters (lines 27-28). During uploading, global rank masks guide parameter pruning (lines 30-31). After receiving client updates, the server aggregates parameters using the FedAvg algorithm [7] (lines 13-14), refines global rank masks via threshold arbitration (lines 15-16), and generates global parameters. Finally, model accuracy is assessed using a test dataset (lines 18-20).
@@ IV Design Details of FedARA
In this section, we present the design of FedARA by elaborating on three interrelated methods: truncated SVD adaptation, dynamic rank allocation, and rank-based module pruning.
@@ IV-A Truncated SVD Adaptation
To address this issue, we draw inspiration from the structure of truncated Singular Value Decomposition (SVD), which decomposes a matrix into the eigenvalues and eigenvectors of its low-rank subspace. Accordingly, we propose a truncated SVD adaptation method that mirrors this structure, formulated as
@@
We now formally provide theoretical insights into the effectiveness of truncated SVD adaptation under non-IID settings. For each client $k$, let the rank–$r$ update be defined as:
@@
and we quantify client heterogeneity by the drift variance:
@@
We adopt the standard separable-correlation model:
@@
where $\rho_{a}\rho_{b}\neq 0$ captures non-IID data. For the truncated SVD adaptation, we further assume:
@@
Assumption (8) is justified since each $e_{i}$ is initialized as $e_{i}^{(0)}=0$ and updated by $e_{i}\leftarrow e_{i}-\eta\langle\nabla\ell(x;W),\,b_{i}a_{i}^{\top}\rangle$, where $e_{i}$ only interacts with its own rank-1 component, leading to independent gradient noise across $i\neq j$. Under symmetric mini-batch sampling, the noise is unbiased with $\mathbb{E}[\Delta e_{i}]=0$. Additionally, with Gaussian initialization, the inner product $(b_{i}^{\top}b_{j})(a_{i}^{\top}a_{j})=O(1/\sqrt{d})$, yielding residual correlation of order $O(1/r)$ when $r<\sqrt{d}$.
Substituting equations (6)–(8), we obtain for LoRA:
$\mathbb{E}\!\left[\|\Delta W^{\text{BA}}\|_{F}^{2}\right]=r\tau_{b}\tau_{a}+r(r-1)\rho_{b}\rho_{a}=\Theta(r^{2}),$
while for truncated SVD adaptation, the $r(r-1)$ cross terms are scaled by $\mathbb{E}[e_{i}e_{j}]$ and vanish up to order $O(r)$, yielding:
$\mathbb{E}\!\left[\|\Delta W^{\text{BEA}}\|_{F}^{2}\right]=r\tau_{e}\tau_{b}\tau_{a}+O(r)=\Theta(r).$
Thus, the variance scales as $\sigma^{2}_{\text{BA}}=\Theta(r^{2})$ and $\sigma^{2}_{\text{BEA}}=\Theta(r)$, confirming that truncated SVD adaptation eliminates the quadratic cross-covariance terms and reduces drift growth from $\Theta(r^{2})$ to $\Theta(r)$. This theoretically supports the effectiveness of the newly diagonal matrix $E$ in mitigating non-IID effects.
@@
This metric measures the averaged cosine similarity between the global and local models, with values closer to 1 indicating higher directional alignment.

### v4

@@ Abstract
[…]
@@ I Introduction
Although PLMs have demonstrated superior performance on general-purpose tasks [3, 15], they still require fine-tuning with local data to achieve optimal results in specific domains. However, since this data is typically generated by users, directly collecting it poses a significant risk of privacy leakage. To address this issue, Federated Learning (FL) is considered an effective approach [25], as it protects data privacy by keeping raw data on local devices, ensuring compliance with regulations such as GDPR [9] and the Data Protection Act [32]. The FL process involves a central server orchestrating the aggregation of local parameter updates from clients.
@@
Figure 1: The complete training process of Federated Parameter-Efficient Fine-Tuning (FedPEFT), where local updates utilize Low-Rank Adaptation (LoRA), while keeping the pre-trained model frozen.
@@
However, the massive scale of PLMs, often with hundreds of millions to billions of parameters [8], incurs substantial computation and communication costs in FL, rendering deployment on resource-constrained devices impossible. To address this issue, a promising solution is to adopt Parameter-Efficient Fine-Tuning (PEFT) methods, which train only a minimal set of parameters using adaptable modules while keeping the pre-trained model frozen [13, 28, 19, 22, 14]. Among these PEFT methods, Low-Rank Adaptation (LoRA) [14] has gained widespread adoption for its efficiency, achieved by introducing two trainable low-rank matrices for diverse task adaptation. By combining PEFT with FL, distributed data resources on local devices can be fully leveraged for collaborative fine-tuning in a privacy-preserving manner. Consequently, Federated Parameter-Efficient Fine-Tuning (FedPEFT) shows significant advancements and finds wide-ranging applications across various domains [40, 34, 2, 1, 37, 38, 31]. Building on this progress, FedPEFT has increasingly gained recognition as a key method for distributed fine-tuning on local devices [4, 35]. Figure 1 shows the complete FedPEFT training process.
@@
Nevertheless, FedPEFT still faces two major challenges when applied to local devices. 1) FedPEFT may exacerbate the impact of non-IID data on model performance. Performance degradation caused by non-independent and identically distributed (non-IID) data, where different clients hold data from distinct distributions, is a well-known challenge in FL. Our preliminary measurements show that FedPEFT exacerbates this issue. Specifically, on the 20News dataset [18] using the DistilBERT model [30], FedPEFT methods exhibit an average accuracy drop of 14.62% from non-IID to IID settings—more than double the 6% drop observed with federated full fine-tuning [1]. 2) Fixed-rank configurations in FedPEFT may incur dual losses in performance and communication cost compared with dynamic-rank allocations. These dual losses arise not because rank choices inherently harm accuracy, but rather because static allocations cannot accommodate the heterogeneous sensitivity of different modules under a constrained total rank budget. In particular, static allocations often underfit high-sensitivity positions, leading to suboptimal performance—our experiments reveal accuracy gaps of up to 9.76% across different configurations. Meanwhile, redundant parameters assigned to less critical components are repeatedly transmitted during FL rounds without corresponding performance gains, resulting in communication inefficiency.
@@
These losses become more severe under non-IID settings, where data heterogeneity intensifies the subspace divergence of federated low-rank adaptation among client-specific models. This divergence leads to greater variability in rank requirements, making fixed-rank configurations increasingly misaligned with actual capacity needs. Consequently, both performance degradation and communication overhead are further amplified. Therefore, these two challenges should be jointly addressed to fully unlock the potential of FedPEFT in practical FL scenarios.
@@
[…]
@@
  • We validate the effectiveness of FedARA through extensive experiments across various datasets and models on advanced edge devices, demonstrating that FedARA consistently improves performance under non-IID data and significantly boosts system efficiency in both communication and computation.
@@ II-A Federated Parameter-Efficient Fine-Tuning (FedPEFT)
FedPEFT orchestrates PEFT within the FL framework to efficiently update local models, drastically reducing communication overhead by only transmitting a small number of PEFT modules. These PEFT modules can be replaced by Adapter [13, 28], Prompt[19, 22], or LoRA [14]. Among them, LoRA-based methods are especially popular because they introduce no additional latency during inference. FedPEFT not only preserves privacy but also effectively aggregates diverse knowledge across clients, facilitating the collaborative development of a high-quality global model. By harnessing the advantages of both PEFT and FL, FedPEFT has become an essential technique for deploying fine-tuning on resource-constrained devices [4, 35, 37, 2, 34, 38].
As illustrated in Figure 1, FedPEFT training process unfolds as follows. Step 1: A cloud server selects a subset of clients to broadcast the initial parameters. Step 2: On these clients, the weights of the pre-trained model are frozen, and trainable PEFT modules are inserted for local training. Step 3: Only these PEFT modules are updated to the cloud server from each client. Step 4: The cloud server aggregates these updates from different clients. This cycle (Steps 1-4) is repeated for multiple rounds in FL until the model converges.
@@ II-B Preliminary Measurements
[…]
@@
Observation 2: The performance of FedPEFT significantly decreases under non-IID data. Figure 2b illustrates the accuracy of different FedPEFT methods using the DistilBERT [30] model for both IID (white bars) and non-IID data (black bars). The non-IID partitioning follows the FedAvg setting [25], with a severe class imbalance in data distribution. The parameter sizes for Adapter-h (A-h) [13] and Adapter-p (A-p) [28] are set to match those of LoRA. Our measurements indicate a significant performance drop in all FedPEFT methods under non-IID conditions compared to IID conditions, with average accuracy reductions of 14.62% and 12.81% on the 20News [18] and News Category [27] datasets, respectively. Existing research lacks effective strategies to mitigate the adverse impact of non-IID data on performance.
@@
Observation 3: The contribution of PEFT modules in FedPEFT to performance varies with their inserted positions, leading to suboptimal performance and inefficient communication in fixed parameter configurations. In FedPEFT, low-rank modules can be placed at various layers and components, such as Query (Q), Key (K), Value (V), Out (O), F1, and F2. Here, O denotes the linear layer in the multi-head attention block, while F1 and F2 represent the two linear layers of the feed-forward network. As shown in Figure 2c, we allocated a fixed rank budget to specific components or layers to evaluate their impact on accuracy. The left sub-figure shows the allocation of the entire rank budget to a single layer, while the right sub-figure illustrates its distribution across a single component in all layers. The findings reveal that deeper layers near the output often have a more significant effect than shallower ones, with components F1 and F2 exerting a greater influence than component O. Moreover, the impact of these inserted positions varies across datasets, suggesting that there is no one-size-fits-all configuration. Fixed-rank configurations result in inefficient communication, as they repeatedly transmit less impactful parameters over rounds, yielding minimal performance gains under a limited parameter budget.
@@
Observation 4: Bottlenecks differ across different types of advanced edge devices. In experiments with LoRA-based methods [40, 34] on News Category [27], given a communication bandwidth of 1MB/s, we found that the bottlenecks varied depending on the device type, as shown in Figure 2d. On high-end devices, communication is the predominant bottleneck, with communication-to-computation ratios of 4.44 and 5.35 for different models. In contrast, on Raspberry Pi devices, computation becomes the primary bottleneck, with communication-to-computation ratios of 0.67 and 0.72. Importantly, the bottleneck shifts towards communication as model parameters increase, consistent with the recent study [35].
@@
Insights: These observations underscore key directions for advancing our research. First, given FedLoRA’s advantages, such as improved training efficiency, future research could benefit from moving beyond comparisons with FedFFT and instead focusing on addressing the unique challenges it presents in FL. Second, the performance degradation of different PEFT methods under non-IID remains a critical barrier, necessitating further investigation into effective approaches. Third, fixed parameter configurations lead to communication inefficiency by repeatedly transmitting less impactful parameters over multiple FL rounds, particularly in resource-constrained environments. Finally, optimizing FedPEFT for edge devices requires further refinement, as different hardware presents distinct bottlenecks. To bridge these gaps, enhancing fine-tuning structures and dynamically adjusting parameters could be essential while simultaneously minimizing both communication and computational overhead.
@@ III Overview of FedARA
Algorithm 1 outlines the process of federated adaptive rank allocation. It starts with parameter initialization and client selection (line 4-6). The server prunes parameters using global rank masks via CommPru (lines 7-8, see Section IV-B3) and broadcasts them to the clients (line 9). Clients perform parallel local training (lines 10-11), updating parameters based on local datasets (lines 21-22). Local rank masks are generated from updated parameters using MaskGen (lines 23-24, see Section IV-B1), determining the number of trainable parameters via RankDet (lines 25-26, see Section IV-C). During uploading, global rank masks guide parameter pruning via CommPru (lines 27-28, see Section IV-B3). After receiving client updates, the server aggregates parameters using the FedAvg algorithm [25] (lines 12-13), refines global rank masks using threshold arbitration via FedArb (lines 14-15, see Section IV-B2), and generates global parameters. Finally, model accuracy is assessed using a test dataset (lines 17-19).
@@ IV Design Details of FedARA
In this section, we present the design of FedARA, consisting of three interrelated methods: truncated SVD adaptation, dynamic rank allocation, and rank-based module pruning.
@@ IV-A Truncated SVD Adaptation
To address this issue, we draw inspiration from the structure of truncated Singular Value Decomposition (SVD), which decomposes a matrix into the singular values and vectors of its low-rank subspace. Accordingly, we propose a truncated SVD adaptation method that mirrors this structure, formulated as
@@
Next, we provide a simplified theoretical analysis to offer intuition for the effectiveness of truncated SVD adaptation. For each client $k$, the rank-$r$ update is defined as:
@@
and we quantify client heterogeneity by the drift variance, assuming zero-mean updates:
@@
To make the analysis tractable, we adopt a simplified separable-covariance assumption:
@@
where $\rho_{a}\rho_{b}\neq 0$, capturing non-trivial cross-rank covariances. This simplified model is inspired by separable (Kronecker-factored) covariance approximations [24]. For the truncated SVD adaptation, we further assume:
@@
Each $e_{i}$ is initialized as $e_{i}^{(0)}=0$ and updated by $e_{i}\leftarrow e_{i}-\eta\langle\nabla\ell(x;W),\,b_{i}a_{i}^{\top}\rangle$, where $e_{i}$ only interacts with its own rank-1 component, leading to independent gradient noise across $i\neq j$. Under symmetric mini-batch sampling, the noise is unbiased with $\mathbb{E}[\Delta e_{i}]=0$. Additionally, with Gaussian initialization, the inner product $(b_{i}^{\top}b_{j})(a_{i}^{\top}a_{j})=O(1/\sqrt{d})$, yielding residual statistical dependence of order $O(1/r)$ when $r<\sqrt{d}$.
Substituting equations (6)-(8), we obtain for LoRA:
$\mathbb{E}\!\left[\|\Delta W^{\text{BA}}\|_{F}^{2}\right]=r\tau_{b}\tau_{a}+r(r-1)\rho_{b}\rho_{a}=\Theta(r^{2}),$ (9)
while for truncated SVD adaptation, the r(r1)r(r-1) cross terms are scaled by $\mathbb{E}[e_{i}e_{j}]$ and vanish up to order $O(r)$, yielding:
$\mathbb{E}\!\left[\|\Delta W^{\text{BEA}}\|_{F}^{2}\right]=r\tau_{e}\tau_{b}\tau_{a}+O(r)=\Theta(r).$ (10)
Thus, the variance scales as $\sigma^{2}_{\text{BA}}=\Theta(r^{2})$ and $\sigma^{2}_{\text{BEA}}=\Theta(r)$, indicating that the diagonal matrix $E$ suppresses quadratic cross-rank covariances induced by non-IID updates. This result is derived under the simplified separable-covariance model. When cross-rank covariances vanish ($\rho_{a}\rho_{b}=0$), all cross terms disappear and both methods yield $\Theta(r)$ variance, leaving little room for further improvement. Accordingly, our analysis focuses on the regime with non-zero cross-rank covariances ($\rho_{a}\rho_{b}\neq 0$) to examine the effect of truncated SVD adaptation, while acknowledging that this simplified model may not strictly hold in practice.
@@
This metric measures the average cosine similarity between the global and local models, with values closer to 1 indicating higher directional alignment.